Two tugboats pull a disabled supertanker. Each tugboat exerts a constant force of 1.8×10⁶, one 14 degrees west of north and the other 14 degrees east of north, as they pull the supertanker 0.751km. Calculate:

A. The total force vector exerted on the supertanker by tugboats as they pull it towards the north.

B. The total work the tugboat do on the supertanker.


1
Expert's answer
2021-02-15T17:38:15-0500


To calculate the total force we should add two vectors using the rule of parallelogram, so we determine the sum of projections of forces on the direction to the north:

"F_{\text{tot}} = 2\cdot F\cdot \cos 14^\circ= 2\cdot 1.8\cdot10^6\,\text{N}\cdot\cos 14^\circ = 3.49\cdot10^6\,\text{N}."


The total work is a product of the total force and the displacement of an object, so

"A = \vec{F}_{\text{tot}} \cdot \vec{r} = {F}_{\text{tot}}\cdot r \cdot\cos0^\circ= 3.49\cdot10^6\,\text{N} \cdot 751\,\mathrm{m} = 2.6\cdot10^9\,\mathrm{J}."